OpenStudy (anonymous):

Including 5% sales tax, an inn charges $210.00 per night. Find the nightly cost at the inn before tax is added. a) Write an equation to model the problem. Let x represent the nightly cost before tax. Answer: b) Solve the equation to find nightly cost before tax. (Note: Your answer should be in the form $ddd.cc.) Answer:

OpenStudy (anonymous):

what is ur question on the question?

OpenStudy (anonymous):

i dong know how to set it up

OpenStudy (anonymous):

ok, Including 5% sales tax, an inn charges $210.00 per night. means 5%*x + x = $210.00

OpenStudy (anonymous):

how do i get the last part. don't i just divide .05 into 210.00

OpenStudy (anonymous):

close! we do want to isolate and solve for x, but we need to do it in the right order. this equation 5%*x + x = $210.00 can be rewritten as 0.5x + x =210 so combine like terms, then divide.

OpenStudy (anonymous):

would it be 1.5x when added together

OpenStudy (anonymous):

yep, so what does x equal?